# Maël Dumas LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Anthony Perez LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Mathis Rocton Algorithms and Complexity Group, TU Wien, Vienna, Austria Ioan Todinca LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France

### v2

@@ 4.2 Reducing Strictly Chordal branches
Let $G=(V,E)$ be a graph. We say that a subgraph $B$ of $G$ is a SC-branch if $V(B)$ is an union of critical cliques $K_{1},\dots,K_{r}\in\mathcal{K}(G)$ such that the subgraph of $\mathcal{C}(G)$ induced by $K_{1},\dots,K_{r}$ is a connected block graph.
We emphasize that our definition of a SC-branch $B$ is stronger than simply requiring $B$ to be an induced strictly chordal graph. For example, if $G$ is the dart graph, the subgraph obtained by removing the pendant vertex is strictly chordal, but it is not a SC-branch because the corresponding critical cliques do not form a block graph in $\mathcal{C}(G)$. Let $B$ be a SC-branch of graph $G$ and let $K_{1},\dots,K_{r}$ be the critical cliques of $G$ contained in $V(B)$. We say that $K_{i}$ is an attachment point of $B$ if $N_{G}(K_{i})\setminus V(B)\neq\emptyset$. A SC-branch $B$ is a $p$-SC-branch if it has exactly $p$ attachment points. We denote $B^{R}$ the subgraph of $B$ in which all attachment points have been removed.
@@
First, we show that all sets $Q_{i}$ are cliques. Suppose that $Q_{i}$ is not a clique for some $1\leqslant i\leqslant r$. Let $x$ and $y$ be non adjacent vertices of $Q_{i}$ and $z\in P$. Since $G_{i}$ is connected, take a shortest path $\pi$ between $x$ and $y$ in $G_{i}$. The subgraph induced by the vertices $\{x,y,z\}$ and those of $\pi$ contains either a cycle of length at least $4$ if $z$ is not adjacent to any inner vertex of $\pi$, which is a contradiction since $B$ is a block graph, or a diamond with $z$ being one of its vertices of degree $3$. In the latter case, since $z$ is not in the same critical clique of $\mathcal{C}(G)$ as its true twin in the diamond, the critical cliques of $\mathcal{C}(G)$ that contain some vertices of this diamond also form a diamond in $\mathcal{C}(G)$. Such a diamond is formed by critical cliques of $G$ contained in $B$, contradicting the definition of a SC-branch. In all cases we have a contradiction, it remains that $Q_{i}$ is a clique in $G$.
Now suppose that $Q_{i}$, $1\leqslant i\leqslant r$ is neither a critical clique nor a maximal clique of $G_{i}$ and intersects at least two maximal cliques of $G_{i}$. Suppose first that $Q_{i}$ is not a module, then there exist $u,v\in Q_{i}$ and $x\in V(G_{i})\setminus Q_{i}$ adjacent to only $u$ or $v$. Since $Q_{i}$ is not a maximal clique of $G_{i}$, there exists $y\in V(G_{i})\setminus Q_{i}$ adjacent to $u$ and $v$. Together with $w\in P$, the vertices $\{u,v,w,x,y\}$ induce in $B$ a gem if $x$ and $y$ are adjacent and a dart otherwise, leading to a contradiction. If $Q_{i}$ is a module, since it is not a critical clique (i.e. a maximal clique module) there exist $u\in Q_{i}$ and $v\in V(G_{i})\setminus Q_{i}$ that lie in the same critical clique of $G_{i}$. Since $Q_{i}$ intersects two maximal cliques of $G_{i}$, there exist two non adjacent vertices $x,y\in V(G_{i})\setminus Q_{i}$ that are adjacent to both $u$ and $v$. Together with $w\in P$, the vertices $\{u,v,w,x,y\}$ induce a dart in $B$, leading to a contradiction. It remains that $Q_{i}$ is a critical clique, a maximal clique or intersects exactly one maximal clique of $G_{i}$. ∎
@@
Suppose that $C^{\prime}$ is not a critical clique of $H^{\prime}$ and intersects at least two maximal cliques of $H^{\prime}$. Since $C^{\prime}\subseteq C$ is a module but not a critical clique, it is not a maximal clique module and thus there exist $u\in C^{\prime}$ and $v\notin C^{\prime}$ that lie in the same critical clique in $H^{\prime}$. Notice that there must exist a vertex $w\in N_{H}(C^{\prime})\cap V(B^{R})$ adjacent to exactly one of $u$ and $v$ in $H$, since otherwise we would have $N_{H}(u)\cap V(B)=N_{H}(v)\cap V(B)$ and $N_{H^{\prime}}(u)=N_{H^{\prime}}(v)$, meaning that $C$ is not a critical clique of $H$, a contradiction. Since $C^{\prime}$ intersects two maximal cliques of $H^{\prime}$ there exist two non adjacent vertices $x,y\in V(H^{\prime})$ that are adjacent to both $u$ and $v$. The vertices $\{u,v,w,x,y\}$ induce a dart in $H$ if $w$ is not adjacent to neither $x$ nor $y$, a gem if it is adjacent to one of them, and if $w$ is adjacent to both $x$ and $y$, $\{w,x,v,y\}$ induce a $C_{4}$ in $H$. It remains that $C^{\prime}$ is a critical clique of $H^{\prime}$ or $C^{\prime}$ is included in exactly one maximal clique of $H^{\prime}$ if $N_{H}(C^{\prime})\cap V(B^{R})\neq\emptyset$. ∎
@@
We now show a result that will allow us to define a rule bounding the number of $1$-SC-branches sharing the same neighborhood. In particular, if enough $1$-SC-branches share a neighborhood $N$, they force any editing of size at most $k$ to transform $N$ into a clique.
@@
We first prove that $N$ has to be clique in $H$. Suppose that $N$ is not a clique in $H$, then there exist two vertices $x,y\in N$ such that $xy\notin E(H)$. The attachment points are critical cliques, so by Rule 4.2, $|P_{i}|\leqslant k+1,i\in\{1,\dots,l\}$. We can observe that there is at most one critical clique $P_{j}$ such that $N_{B_{j}}(P_{j})=\emptyset$ (recall that $r=1$ and hence all attachment points lie in the same clique in $G[P_{1}\cup\dots\cup P_{l}]$). It follows that $|Q_{1}\setminus P_{j}|>k$, thus there exists a vertex $u$ in some $P_{t},t\neq j$ such that in $H$, $u$ is adjacent to a vertex $z\in N_{B_{t}}(P_{t})$ and the vertices $x$ and $y$. Since $|Q_{1}\setminus P_{t}|>k$, there exists a vertex $v\in Q_{1}\setminus P_{t}$ such that there is no pair of $F$ containing both $v$ and a vertex of $\{x,y,z,u\}$. The set $\{x,y,u,v,z\}$ induces a dart if neither $x$ nor $y$ is adjacent to $z$, a gem if one of them is adjacent to $z$, or if both of them are adjacent to $z$, $\{x,y,z,v\}$ induces a $C_{4}$ in $H$, a contradiction. It follows that $N$ has to be a clique.
We now show that $N$ is a critical clique, a maximal clique or intersects exactly one maximal clique of $H^{\prime}$. Suppose that $N$ is neither a critical clique nor a maximal clique of $H^{\prime}$ and intersects at least two maximal cliques of $H^{\prime}$. Suppose first that $N$ is not a module, there exist $u,v\in N$, $x\in V(H^{\prime})\setminus N$ adjacent to only $u$ or $v$. Since $N$ is not a maximal clique of $H^{\prime}$, there exist $y\in V(H^{\prime})\setminus N$ adjacent to $u$ and $v$. Together with a vertex $w\in\bigcup_{i=1}^{l}P_{i}$ not affected by $F$, the vertices $\{u,v,w,x,y\}$ induce in $H$ a gem if $x$ and $y$ are adjacent and a dart else, leading to a contradiction. If $N$ is a module and is not a critical clique, $N$ is not a maximal clique module, thus there exist $u\in N$, a vertex $v$ in the same module as $u$ in $H^{\prime}$ and a vertex $w\in\bigcup_{i=1}^{l}P_{i}$ not affected by $F$ such that $vw\notin E(H)$. Since $N$ intersects two maximal cliques of $H^{\prime}$ there exist two non adjacent vertices $x,y\in V(H^{\prime})$ that are adjacent to both $u$ and $v$. The vertices $\{u,v,w,x,y\}$ induce a dart in $H$ since $w$ is only adjacent to $u$. It remains that if $r=1$, $N$ is a critical clique, a maximal clique or is included in exactly one maximal clique of $H^{\prime}$.
@@
Finally, since $N$ is a module in $H^{\prime}$, if it is not a critical clique, there exist $u\in N$, a vertex $v$ in the same critical clique as $u$ in $H^{\prime}$ and a vertex $w\in\bigcup_{i=1}^{l}P_{i}$ not affected by $F$ such that $vw\notin E(H)$. If $N$ intersects at least two maximal cliques of $H^{\prime}$ there exist two non adjacent vertices $x,y\in V(H^{\prime})$ that are adjacent to both $u$ and $v$. The vertices $\{u,v,w,x,y\}$ induce a dart in $H$ since $w$ is only adjacent to $u$. It remains that $N$ is a critical clique or intersects exactly one maximal clique of $H^{\prime}$. ∎
@@ 4.3 Reducing the 2-SC-branches
Assume now that $P_{1}$ and $P_{2}$ are in the same connected component in $H^{\prime}$ and let $\pi_{H^{\prime}}$ denote a shortest path between $c_{1}\in P_{1}$ and $c_{2}\in P_{2}$. Note that its length is potentially $1$ if $c_{1}$ and $c_{2}$ are adjacent in $H^{\prime}$. We first consider the case where there is a path $\pi_{B}=\{c_{1}=u_{1},u_{2},\dots,u_{r}=c_{2}\}$ (not necessarily induced) from $c_{1}$ to $c_{2}$ in $B$ that still exists in $B\triangle F$. Observe that such this path $\pi_{B}$ is of length at least $k+3$ since $B$ is of length at least $k+3$. There is no edge between the two paths in $G$, so $H[V(\pi_{B})\cup V(\pi_{H})]$ admits at least one induced cycle of length at least 4, which is a contradiction. We conclude that if $P_{1}$ and $P_{2}$ are in the same connected component of $H\setminus V(B^{R})$, then $F$ contains an edge-cut of $B$. Else there would be a path $\pi_{B}$ from $c_{1}$ to $c_{2}$ in $B$ that still exists in $B\triangle F$.
@@
Now, if $F$ contains neither $(P_{1}\times N_{B}(P_{1}))$ nor $(P_{2}\times N_{B}(P_{2}))$, let $F_{1}=F\cap(V(B^{R})\times V(B^{R}))$. There are two components in $B\triangle F_{1}$, say $B_{1}$ the one containing $P_{1}$ and $B_{2}$ one containing $P_{2}$. These connected components are $1$-SC-branches of $G_{1}=G\triangle F_{1}$ with attachment points $P_{1}$ and $P_{2}$. As before Lemma 17 applies on the $1$-SC-branches $B_{1}$ and $B_{2}$, thus there is an optimal edition $F_{2}$ of $G_{1}$ where the only vertices of $B_{1}$ and $B_{2}$ affected are included in $P_{1}\cup N_{B}(P_{1})$ and $P_{2}\cup N_{B}(P_{2})$. If $F_{1}$ is a min-cut of $B$, the edition $F_{1}\cup F_{2}$ is an optimal edition of $G$ that verifies the desired properties, else, replacing $F_{1}$ by a min-cut of $B$ result in a smaller edition that verifies the properties. ∎
@@ 4.4 Bounding the size of the kernel
Lemma 24 ([36]).
Let $G=(V,E)$ be a graph and $H=G\triangle\{uv\}$ for $u,v\in V$, then $|\mathcal{K}(H)|\leqslant|\mathcal{K}(G)|+2$.
Theorem 4.
Strictly Chordal Editing admits a kernel with $O(k^{4})$ vertices.
@@
Let $(G=(V,E),k)$ be a reduced yes-instance of Strictly Chordal Editing, $F$ a $k$-édition of $G$ and $H=G\triangle F$. We assume that $G$ is connected, the following arguments can be easily adapted if this is not the case by summing over all connected components of $G$. The graph $H$ is strictly chordal, thus the graph $\mathcal{C}(H)$ is a block graph. We first show that $\mathcal{C}(H)$ has $O(k^{3})$ vertices. We say that a critical clique of $\mathcal{C}(H)$ is affected if it contains a vertex affected by $F$. Let $A$ be the set of affected critical cliques of $\mathcal{C}(H)$. Since $|F|\leqslant k$, we have $|A|\leqslant 2k$. Let $T$ be the minimal connected induced subgraph of $H$ that spans all critical cliques of $A$, and $A^{\prime}$ the set of vertices of degree at least $3$ in $T$ (see Figure 8). Recall that such a subgraph is unique by Lemma 5.
@@
We first show that $|V(T)|$ contains $O(k^{2})$ critical cliques. By Lemma 5, $|A^{\prime}\setminus A|\leqslant 3\cdot|A|\leqslant 6k$. The connected components of the graph $T\setminus(A\cup A^{\prime})$ are paths since every vertex is of degree at most $2$ and by Lemma 5 there are at most $4k$ such paths. Let $R$ be one of these paths, it is composed of unaffected critical cliques, thus there exists a clean $2$-SC-branch $B$ of $G$ that contains $R$. Moreover, the extremities of $R$ are the attachment points of $B$, hence $B$ have been reduced by Rule 4.5. Thus $R$ is of length at most $k+3$. It remains that $|V(T)|\leqslant 2k+6k+(k+3)\cdot 4k=O(k^{2})$.
@@
We will now show that $\mathcal{C}(H)\setminus V(T)$ contains $O(k^{3})$ critical cliques. First observe that each connected component of $\mathcal{C}(H)\setminus V(T)$ is adjacent to some vertices of $T$ since the graph is reduced by Rule 4.1. Since $\mathcal{C}(H)$ is a block graph, connected components of $\mathcal{C}(H)\setminus V(T)$ are adjacent to either a critical clique of $T$ or a maximal clique of $T$ (else, there would be a diamond in $\mathcal{C}(H)$). We claim that there are $O(k^{3})$ critical cliques in the connected components of $\mathcal{C}(H)\setminus V(T)$ adjacent to maximal cliques of $T$. Since $T$ is a block graph and $|V(T)|=O(k^{2})$, there are $O(k^{2})$ maximal cliques in $T$. Moreover, there is only one connected component of $\mathcal{C}(H)\setminus V(T)$ adjacent to each maximal clique, otherwise there would be a diamond in $\mathcal{C}(H)$. Take $K$ a maximal clique of $T$ and let $X$ be its adjacent connected component of $\mathcal{C}(H)\setminus V(T)$. Observe that $X$ has to be an union of $1$-SC-branches and their attachment points form a clique. By Rule 4.4, there are at most $12k+$ 1 $1$-SC-branches in $X$ and each one has been reduced by Rule 4.3, thus $X$ contains at most $24k+$ 2 critical cliques.
@@
Finally, we claim that there are $O(k^{2})$ critical cliques in the connected components of $\mathcal{C}(H)\setminus V(T)$ adjacent to critical cliques of $T$. First take a critical clique of $T\setminus A$ and its adjacent connected components of $\mathcal{C}(H)\setminus V(T)$, recall that $|T\setminus A|=O(k^{2})$. Observe that they form a $1$-SC-branch, reduced by Rule 4.3, thus the adjacent connected component actually consists in only one critical clique. Next, take a critical clique $a$ of $A$, and let $C_{1},\dots,C_{r}$ be the connected components of $\mathcal{C}(H)\setminus V(T)$ adjacent to $a$ and $Y$ the union of these connected components (see Figure 8). Observe that each $C_{i}$ has to be an union of $1$-SC-branches and their attachment points form a clique $Q_{i}$. Let $B_{1},\dots,B_{l}$ with attachment points $P_{1},\dots,P_{l}$ be the $1$-SC-branches of $Y$. By Rule 4.4, there are at most $12k+$ 1 $1$-SC-branches by connected component and if $\Sigma_{i=1}^{l}|P_{i}|>2k+1$, then $(\Sigma_{i=1}^{l}|P_{i}|)-\max_{1\leqslant j\leqslant r}\{|Q_{j}|\}\leqslant k$, implying that there are at most $13k+$ 1 $1$-SC-branches in $Y$. Each of these $1$-SC-branches is reduced by Rule 4.3, hence $Y$ contains at most $26k+$ 2 critical cliques in total. Since $|A|=O(k)$, there is $O(k)$ such sets of $1$-SC-branches in $\mathcal{C}(H)\setminus V(T)$. Altogether, it remains that $\mathcal{C}(H)\setminus V(T)$ contains $O(k^{2})$ critical cliques adjacent to critical cliques of $T$.
@@
Figure 8: Illustration of the critical clique graph of a reduced instance in the proof of Theorem 4. Square nodes correspond to vertices of $T$, the ones filled in red are vertices of $A$, the ones filled in blue are vertices of $A^{\prime}\setminus A$.

### v3

@@ 4.2 Reducing Strictly Chordal branches
Let $G=(V,E)$ be a graph. We say that a subgraph $B$ of $G$ is a SC-branch if $V(B)$ is a union of critical cliques $K_{1},\dots,K_{r}\in\mathcal{K}(G)$ such that the subgraph of $\mathcal{C}(G)$ induced by $K_{1},\dots,K_{r}$ is a connected block graph.
We emphasize that our definition of a SC-branch $B$ is stronger than simply requiring $B$ to be an induced strictly chordal graph. For example, if $G$ is the dart graph, the subgraph obtained by removing the pendant vertex is strictly chordal, but it is not a SC-branch because the corresponding critical cliques do not form a block graph in $\mathcal{C}(G)$. Let $B$ be a SC-branch of graph $G$ and let $K_{1},\dots,K_{r}$ be the critical cliques of $G$ contained in $V(B)$. We say that $K_{i}$ is an attachment point of $B$ if $N_{G}(K_{i})\setminus V(B)\neq\emptyset$. A SC-branch $B$ is a $p$-SC-branch if it has exactly $p$ attachment points. We denote $B^{R}$ the induced subgraph of $B$ in which all attachment points have been removed.
@@
First, we show that all sets $Q_{i}$ are cliques. Suppose that $Q_{i}$ is not a clique for some $1\leqslant i\leqslant r$. Let $x$ and $y$ be non-adjacent vertices of $Q_{i}$ and $z\in P$. Since $G_{i}$ is connected, take a shortest path $\pi$ between $x$ and $y$ in $G_{i}$. The subgraph induced by the vertices $\{x,y,z\}$ and those of $\pi$ contains either a cycle of length at least $4$ if $z$ is not adjacent to any inner vertex of $\pi$, which is a contradiction since $B$ is a block graph, or a diamond with $z$ being one of its vertices of degree $3$. In the latter case, since $z$ is not in the same critical clique of $\mathcal{C}(G)$ as its true twin in the diamond, the critical cliques of $\mathcal{C}(G)$ that contain some vertices of this diamond also form a diamond in $\mathcal{C}(G)$. Such a diamond is formed by critical cliques of $G$ contained in $B$, contradicting the definition of a SC-branch. In all cases we have a contradiction, it remains that $Q_{i}$ is a clique in $G$.
Now suppose that $Q_{i}$, $1\leqslant i\leqslant r$ is neither a critical clique nor a maximal clique of $G_{i}$ and intersects at least two maximal cliques of $G_{i}$. Suppose first that $Q_{i}$ is not a module, then there exist $u,v\in Q_{i}$ and $x\in V(G_{i})\setminus Q_{i}$ adjacent to only $u$ or $v$. Since $Q_{i}$ is not a maximal clique of $G_{i}$, there exists $y\in V(G_{i})\setminus Q_{i}$ adjacent to $u$ and $v$. Together with $w\in P$, the vertex set $\{u,v,w,x,y\}$ induces in $B$ a gem if $x$ and $y$ are adjacent and a dart otherwise, leading to a contradiction. If $Q_{i}$ is a module, since it is not a critical clique (i.e. a maximal clique module) there exist $u\in Q_{i}$ and $v\in V(G_{i})\setminus Q_{i}$ that lie in the same critical clique of $G_{i}$. Since $Q_{i}$ intersects two maximal cliques of $G_{i}$, there exist two non-adjacent vertices $x,y\in V(G_{i})\setminus Q_{i}$ that are adjacent to both $u$ and $v$. Together with $w\in P$, the vertex set $\{u,v,w,x,y\}$ induces a dart in $B$, leading to a contradiction. It remains that $Q_{i}$ is a critical clique, a maximal clique or intersects exactly one maximal clique of $G_{i}$. ∎
@@
Suppose that $C^{\prime}$ is not a critical clique of $H^{\prime}$ and intersects at least two maximal cliques of $H^{\prime}$. Since $C^{\prime}\subseteq C$ is a module but not a critical clique, it is not a maximal clique module and thus there exist $u\in C^{\prime}$ and $v\notin C^{\prime}$ that lie in the same critical clique in $H^{\prime}$. Notice that there must exists a vertex $w\in N_{H}(C^{\prime})\cap V(B^{R})$ adjacent to exactly one of $u$ and $v$ in $H$, since otherwise we would have $N_{H}(u)\cap V(B)=N_{H}(v)\cap V(B)$ and $N_{H^{\prime}}(u)=N_{H^{\prime}}(v)$, meaning that $C$ is not a critical clique of $H$, a contradiction. Since $C^{\prime}$ intersects two maximal cliques of $H^{\prime}$ there exist two non-adjacent vertices $x,y\in V(H^{\prime})$ that are adjacent to both $u$ and $v$. The vertex set $\{u,v,w,x,y\}$ induces a dart in $H$ if $w$ is not adjacent to neither $x$ nor $y$, a gem if it is adjacent to one of them, and if $w$ is adjacent to both $x$ and $y$, $\{w,x,v,y\}$ induces a $C_{4}$ in $H$. It remains that $C^{\prime}$ is a critical clique of $H^{\prime}$ or $C^{\prime}$ is included in exactly one maximal clique of $H^{\prime}$ if $N_{H}(C^{\prime})\cap V(B^{R})\neq\emptyset$. ∎
@@
We now show a result that will allow us to define a rule bounding the number of $1$-SC-branches sharing the same neighborhood. In particular, if enough $1$-SC-branches share a neighborhood $N$, they force any edition of size at most $k$ to transform $N$ into a clique.
@@
We first prove that $N$ has to be clique in $H$. Suppose that $N$ is not a clique in $H$, then there exist two vertices $x,y\in N$ such that $xy\notin E(H)$. The attachment points are critical cliques, so by Rule 4.2, $|P_{i}|\leqslant k+1,i\in\{1,\dots,l\}$. We can observe that there is at most one critical clique $P_{j}$ such that $N_{B_{j}}(P_{j})=\emptyset$. Indeed, since $r=1$ all attachment points lie in the same clique in $G[P_{1}\cup\dots\cup P_{l}]$, we have that $N_{G}[P_{i}]\setminus N_{B_{i}}(P_{i})=N\cup P_{1}\cup\dots\cup P_{l}$, hence there is only one such $P_{j}$. It follows that $|Q_{1}\setminus P_{j}|>k$, thus there exists a vertex $u$ in some $P_{t},t\neq j$ such that in $H$, $u$ is adjacent to a vertex $z\in N_{B_{t}}(P_{t})$ and the vertices $x$ and $y$. Since $|Q_{1}\setminus P_{t}|>k$, there exists a vertex $v\in Q_{1}\setminus P_{t}$ such that there is no pair of $F$ containing both $v$ and a vertex of $\{x,y,z,u\}$. The set $\{x,y,u,v,z\}$ induces a dart if neither $x$ nor $y$ is adjacent to $z$, a gem if one of them is adjacent to $z$, or if both of them are adjacent to $z$, $\{x,y,z,v\}$ induces a $C_{4}$ in $H$, a contradiction. It follows that $N$ has to be a clique.
We now show that $N$ is a critical clique, a maximal clique or intersects exactly one maximal clique of $H^{\prime}$. Suppose that $N$ is neither a critical clique nor a maximal clique of $H^{\prime}$ and intersects at least two maximal cliques of $H^{\prime}$. Suppose first that $N$ is not a module, there exist $u,v\in N$, $x\in V(H^{\prime})\setminus N$ adjacent to only $u$ or $v$. Since $N$ is not a maximal clique of $H^{\prime}$, there exists $y\in V(H^{\prime})\setminus N$ adjacent to $u$ and $v$. Together with a vertex $w\in\bigcup_{i=1}^{l}P_{i}$ not affected by $F$, the vertex set $\{u,v,w,x,y\}$ induces in $H$ a gem if $x$ and $y$ are adjacent and a dart else, leading to a contradiction. If $N$ is a module and is not a critical clique, $N$ is not a maximal clique module, thus there exist $u\in N$, a vertex $v$ in the same module as $u$ in $H^{\prime}$ and a vertex $w\in\bigcup_{i=1}^{l}P_{i}$ not affected by $F$ such that $vw\notin E(H)$. Since $N$ intersects two maximal cliques of $H^{\prime}$ there exist two non-adjacent vertices $x,y\in V(H^{\prime})$ that are adjacent to both $u$ and $v$. The vertex set $\{u,v,w,x,y\}$ induces a dart in $H$ since $w$ is only adjacent to $u$. It remains that if $r=1$, $N$ is a critical clique, a maximal clique or is included in exactly one maximal clique of $H^{\prime}$.
@@
Finally, since $N$ is a module in $H^{\prime}$, if it is not a critical clique, there exist $u\in N$, a vertex $v$ in the same critical clique as $u$ in $H^{\prime}$ and a vertex $w\in\bigcup_{i=1}^{l}P_{i}$ not affected by $F$ such that $vw\notin E(H)$. If $N$ intersects at least two maximal cliques of $H^{\prime}$ there exist two non-adjacent vertices $x,y\in V(H^{\prime})$ that are adjacent to both $u$ and $v$. The vertex set $\{u,v,w,x,y\}$ induces a dart in $H$ since $w$ is only adjacent to $u$. It remains that $N$ is a critical clique or intersects exactly one maximal clique of $H^{\prime}$. ∎
@@ 4.3 Reducing the 2-SC-branches
Assume now that $P_{1}$ and $P_{2}$ are in the same connected component in $H^{\prime}$ and let $\pi_{H^{\prime}}$ denote a shortest path between $c_{1}\in P_{1}$ and $c_{2}\in P_{2}$. Note that $\pi_{H^{\prime}}$ may be of length $1$ if $c_{1}$ and $c_{2}$ are adjacent in $H^{\prime}$. We first consider the case where there is a path $\pi_{B}=\{c_{1}=u_{1},u_{2},\dots,u_{r}=c_{2}\}$ (not necessarily induced) from $c_{1}$ to $c_{2}$ in $B$ that still exists in $B\triangle F$. Observe that such this path $\pi_{B}$ is of length at least $k+3$ since $B$ is of length at least $k+3$. There is no edge between the two paths in $G$, so $H[V(\pi_{B})\cup V(\pi_{H})]$ admits at least one induced cycle of length at least 4, which is a contradiction. We conclude that if $P_{1}$ and $P_{2}$ are in the same connected component of $H\setminus V(B^{R})$, then $F$ contains an edge-cut of $B$. Else there would be a path $\pi_{B}$ from $c_{1}$ to $c_{2}$ in $B$ that still exists in $B\triangle F$.
@@
Now, if $F$ contains neither $(P_{1}\times N_{B}(P_{1}))$ nor $(P_{2}\times N_{B}(P_{2}))$, let $F_{1}=F\cap(V(B^{R})\times V(B^{R}))$. There are two components in $B\triangle F_{1}$, say $B_{1}$ the one containing $P_{1}$ and $B_{2}$ one containing $P_{2}$. These connected components are $1$-SC-branches of $G_{1}=G\triangle F_{1}$ with attachment points $P_{1}$ and $P_{2}$. As before Lemma 17 applies on the $1$-SC-branches $B_{1}$ and $B_{2}$, thus there is an optimal edition $F_{2}$ of $G_{1}$ where the only vertices of $B_{1}$ and $B_{2}$ affected are included in $P_{1}\cup N_{B}(P_{1})$ and $P_{2}\cup N_{B}(P_{2})$. If $F_{1}$ is a min-cut of $B$, the edition $F_{1}\cup F_{2}$ is an optimal edition of $G$ that verifies the desired properties, else, replacing $F_{1}$ by a min-cut of $B$ results in a smaller edition that verifies the properties. ∎
@@ 4.4 Bounding the size of the kernel
See 2
@@
Let $(G=(V,E),k)$ be a reduced yes-instance of Strictly Chordal Editing, $F$ a $k$-edition of $G$ and $H=G\triangle F$. We assume that $G$ is connected, the following arguments can be easily adapted if this is not the case by summing over all connected components of $G$. The graph $H$ is strictly chordal, thus the graph $\mathcal{C}(H)$ is a block graph. We first show that $\mathcal{C}(H)$ has $O(k^{3})$ vertices. We say that a critical clique of $\mathcal{C}(H)$ is affected if it contains a vertex affected by $F$. Let $A$ be the set of affected critical cliques of $\mathcal{C}(H)$. Since $|F|\leqslant k$, we have $|A|\leqslant 2k$. Let $T_{A}$ be the minimal connected induced subgraph of $H$ that spans all critical cliques of $A$, and $d(T_{A})$ the set of vertices of degree at least $3$ in $T_{A}$ (see Figure 9). Recall that such a subgraph is unique by Lemma 5.
@@
We first show that $|V(T_{A})|$ contains $O(k^{2})$ critical cliques. By Lemma 5, $|d(T_{A})\setminus A|\leqslant 3\cdot|A|\leqslant 6k$. The connected components of the graph $T_{A}\setminus(A\cup d(T_{A}))$ are paths since every vertex is of degree at most $2$ and by Lemma 5 there are at most $4k$ such paths. Let $R$ be one of these paths, it is composed of unaffected critical cliques, thus there exists a clean $2$-SC-branch $B$ of $G$ that contains $R$. Moreover, the endpoints of $R$ are the attachment points of $B$, hence $B$ have been reduced by Rule 4.5. Thus $R$ is of length at most $k+3$. It remains that $|V(T_{A})|\leqslant 2k+6k+(k+3)\cdot 4k=O(k^{2})$.
@@
We will now show that $\mathcal{C}(H)\setminus V(T_{A})$ contains $O(k^{3})$ critical cliques. First observe that each connected component of $\mathcal{C}(H)\setminus V(T_{A})$ is adjacent to some vertices of $T_{A}$ since the graph is reduced by Rule 4.1. Since $\mathcal{C}(H)$ is a block graph, connected components of $\mathcal{C}(H)\setminus V(T_{A})$ are adjacent to either a critical clique of $T_{A}$ or a maximal clique of $T_{A}$ (else, there would be a diamond in $\mathcal{C}(H)$). We claim that there are $O(k^{3})$ critical cliques in the connected components of $\mathcal{C}(H)\setminus V(T_{A})$ adjacent to maximal cliques of $T_{A}$. Since $T_{A}$ is a block graph and $|V(T_{A})|=O(k^{2})$, there are $O(k^{2})$ maximal cliques in $T_{A}$. Moreover, there is only one connected component of $\mathcal{C}(H)\setminus V(T_{A})$ adjacent to each maximal clique, otherwise there would be a diamond in $\mathcal{C}(H)$. Take $K$ a maximal clique of $T_{A}$ and let $X$ be its adjacent connected component of $\mathcal{C}(H)\setminus V(T_{A})$. Observe that $X$ has to be a union of $1$-SC-branches and their attachment points form a clique. By Rule 4.4, there are at most $12k+$ 1 $1$-SC-branches in $X$ and each one has been reduced by Rule 4.3, thus $X$ contains at most $24k+$ 2 critical cliques.
@@
Finally, we claim that there are $O(k^{2})$ critical cliques in the connected components of $\mathcal{C}(H)\setminus V(T_{A})$ adjacent to critical cliques of $T_{A}$. First take a critical clique of $T_{A}\setminus A$ and its adjacent connected components of $\mathcal{C}(H)\setminus V(T_{A})$, recall that $|T_{A}\setminus A|=O(k^{2})$. Observe that they form a $1$-SC-branch, reduced by Rule 4.3, thus the adjacent connected component actually consists in only one critical clique. Next, take a critical clique $a$ of $A$, and let $C_{1},\dots,C_{r}$ be the connected components of $\mathcal{C}(H)\setminus V(T_{A})$ adjacent to $a$ and $Y$ the union of these connected components (see Figure 9). Observe that each $C_{i}$ has to be a union of $1$-SC-branches and their attachment points form a clique $Q_{i}$. Let $B_{1},\dots,B_{l}$ with attachment points $P_{1},\dots,P_{l}$ be the $1$-SC-branches of $Y$. By Rule 4.4, there are at most $12k+$ 1 $1$-SC-branches by connected component and if $\Sigma_{i=1}^{l}|P_{i}|>2k+1$, then $(\Sigma_{i=1}^{l}|P_{i}|)-\max_{1\leqslant j\leqslant r}\{|Q_{j}|\}\leqslant k$, implying that there are at most $13k+$ 1 $1$-SC-branches in $Y$. Each of these $1$-SC-branches is reduced by Rule 4.3, hence $Y$ contains at most $26k+$ 2 critical cliques in total. Since $|A|=O(k)$, there is $O(k)$ such sets of $1$-SC-branches in $\mathcal{C}(H)\setminus V(T_{A})$. Altogether, it remains that $\mathcal{C}(H)\setminus V(T_{A})$ contains $O(k^{2})$ critical cliques adjacent to critical cliques of $T_{A}$.
@@
Figure 9: Illustration of the critical clique graph of a reduced instance in the proof of Theorem 2. Square nodes correspond to vertices of $T_{A}$, the ones filled in red are vertices of $A$, the ones filled in blue are vertices of $d(T_{A})\setminus A$.